# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v1

@@ Abstract
Music recordings often suffer from audio quality issues such as excessive reverberation, distortion, clipping, tonal imbalances, and a narrowed stereo image, especially when created in non-professional settings without specialized equipment or expertise. These problems are typically corrected using separate specialized tools and manual adjustments. In this paper, we introduce SonicMaster, the first unified generative model for music restoration and mastering that addresses a broad spectrum of audio artifacts with text-based control. SonicMaster is conditioned on natural language instructions to apply targeted enhancements, or can operate in an automatic mode for general restoration. To train this model, we construct the SonicMaster dataset, a large dataset of paired degraded and high-quality tracks by simulating common degradation types with nineteen degradation functions belonging to five enhancements groups: equalization, dynamics, reverb, amplitude, and stereo. Our approach leverages a flow-matching generative training paradigm to learn an audio transformation that maps degraded inputs to their cleaned, mastered versions guided by text prompts. Objective audio quality metrics demonstrate that SonicMaster significantly improves sound quality across all artifact categories. Furthermore, subjective listening tests confirm that listeners prefer SonicMaster’s enhanced outputs over the original degraded audio, highlighting the effectiveness of our unified approach. Samples, code, model, dataset, and demo are all available through https://amaai-lab.github.io/SonicMaster/
Introduction
@@
Music tracks produced in amateur settings often suffer from a variety of audio quality issues that distinguish them from professionally mastered recordings (Wilson and Fazenda 2016; Mourgela et al. 2024; Deruty and Tardieu 2014). For instance, an enthusiast recording vocals in a garage may introduce excessive reverberation, making the voice sound distant and “echoey.” Similarly, using inexpensive microphones or misconfigured interfaces can lead to distortion and clipping when loud peaks exceed the recording range, resulting in harsh crackles or flattened dynamics (Zang et al. 2025). Tonal imbalances are also common: a home recording might sound overly “muddy” or “tinny” if certain frequency bands dominate or vanish due to poor room acoustics or improper microphone placement. Even the stereo image can be narrowed or skewed, reducing the sense of space in the mix. In practice, engineers address these problems with separate specialized tools: e.g., dereverberation plugins to remove room echo, declipping algorithms to reconstruct saturated peaks, equalizers to rebalance frequencies, and stereo enhancers to widen the image. Mastering a flawed track thus becomes a labor-intensive process requiring expert skill and multiple stages of manual adjustment.
@@
The need for an automated all-in-one solution is evident. Creators with limited resources often lack the expertise to apply the right combination of restoration tools, and a piecemeal approach may fail to fully recover a track’s fidelity. This motivates SonicMaster, a unified approach to music restoration and mastering that can correct a broad spectrum of audio degradations within a single model.
Figure 1: SonicMaster pipeline: a flow-matching model trained on the SonicMaster dataset (19 simulated enhancements across EQ, dynamics, reverb, amplitude, stereo) takes an input music waveform and text prompts to produce a high-fidelity enhanced output.
@@
We introduce a single flow-based generative framework that simultaneously performs dereverberation, equalization, declipping, dynamic-range expansion, and stereo enhancement. The backbone is trained on a curated corpus of polyphonic music rendered through a combinatorial grid of simulated degradations, enabling the network to learn the joint statistics and cross-couplings of common artifacts rather than treating them in isolation. This joint training eliminates the need for error-prone cascades of task-specific modules and reduces inference to a single forward pass.
SonicMaster is conditioned on free-form text prompts that encode high-level production intent. A prompt such as reduce the hollow room sound attenuates late reflections without suppressing desirable early reverberation, whereas increase the brightness selectively enhances the treble frequencies while while preserving spectral balance elsewhere.When the prompt is omitted, an automatic mode leverages learned perceptual heuristics to produce a balanced master suited to non-expert users. Existing models (e.g. VoiceFixer (Liu et al. 2021)) fail on music-focused pipelines typically address artifacts sequentially, ignoring their mutual influence. By unifying restoration and mastering tasks under a single, prompt-driven generative model, SonicMaster delivers professional-grade improvements while affording fine-grained creative control.
In the absence of text-conditioned music-restoration data, we build a new large-scale corpus for controllable restoration. From $\approx 580$ k Jamendo recordings we retain $\approx 25$ k high-quality 30-s segments, balanced across 10 genre groups by production quality score. Each clean clip is corrupted with one to three of 19 common effects drawn from five categories—EQ, dynamics, reverb, amplitude, and stereo—producing paired degraded versions. Every degraded sample is accompanied by a natural-language prompt describing the artifact or required fix, and all random effect parameters are stored as metadata. This genre-diverse collection of tens of thousands of prompt–audio pairs underpins SonicMaster training and offers a rigorous benchmark for controllable music-restoration research. Our main contributions are as follows:
  • We present SonicMaster, a flow-matching model that corrects $19$ common degradations—including reverb, EQ imbalance, clipping, dynamic-range errors, and stereo artifacts in a single generative pass, while allowing free-form textual prompts for fine-grained creative control.
  • To train and evaluate such a system, we curate and release the first text-conditioned music-restoration corpus: 25k high-fidelity Jamendo segments spanning 10 genres, each paired with 7 degraded versions of itself, detailed parameter metadata, and a natural-language instruction describing the required fix, resulting in 175 k audio pairs.
Related Work
Restoring and mastering audio spans speech + music enhancement, audio inpainting, and source separation—areas that have mostly been handled separately (Záviška et al. 2020). Diffusion-based generative models and text-guided audio editing (Hou et al. 2025; Jiang et al. 2025) now let us tackle these problems together. We review these advances, their uses, and the gaps that SonicMaster aims to fill.
Audio Inpainting and Declipping:
Early signal-model and interpolation methods could patch only very short gaps (<10<10 ms), leaving longer dropouts unresolved. Deep generative models now bridge that gap: diffusion-based systems convincingly regenerate missing music sections and clipped peaks (Moliner and Välimäki 2023). The authors in (Wang et al. 2023; Liu et al. 2023) extend this with instruction-guided diffusion for audio inpainting, while VoiceFixer (Liu et al. 2021) jointly denoises, dereverbs, and declipse speech, though it is restricted to voice and does not offer user control. In music, (Imort et al. 2022) removed heavy guitar distortion (including clipping) with neural networks, surpassing sparse-optimization baselines in quality and speed.
Equalization and Tonal Restoration:
Research on learning-based equalization is still emerging. In recently proposed work (Mockenhaupt, Rieber, and Nercessian 2024), the authors introduce CNN based approach to automatically equalize instrument stems by predicting parametric EQ settings, showing improvements over earlier heuristic methods. Notably, the VoiceFixer (Liu et al. 2021) addressed bandwidth extension essentially restoring high-frequency content as part of its speech restoration, which can be seen as a form of equalization correction. Similarly, diffusion-based restorers like MaskSR (Li, Wang, and Liu 2024) treat low-frequency muffling as a distortion to fix, using discrete token prediction to restore a balanced spectrum. In Text2FX (Chu et al. 2025), CLAP (Elizalde et al. 2023) is used in inference mode to steer the parameters of EQ and reverb audio effects.
Method
Dataset
Due to the absence of natural language instruction-prompted music restoration datasets, we have generated SonicMaster music restoration and mastering dataset with text prompts. We source 580580k recordings from the Jamendo [^1] https://www.jamendo.com/ under creative commons licence using the official Jamendo API. To ensure equal representation of genre, we defined $10$ groups, where each group consists of multiple related genre tags, e.g., Hip-Hop genre group containing the following tags: ”rap”, ”hiphop”, ”trap”, ”alternativehiphop”, ”gangstarap”. The full list can be viewed in Appendix. To pick songs for each genre group, we started by deploying the Audiobox Aesthetics toolbox (Tjandra et al. 2025) to assess the production quality of all the recordings. Then, we selected $2,500$ songs for each genre group by filtering songs based on the production quality with an adaptive threshold between 6.586.5-8, depending on the current genre group. We strived to strike a balance between representing all the subtags in each genre group sufficiently enough while maintaining high enough production quality scores. Finally, we extracted a random 30 sec excerpt from each track, positioned between 1515% and 8585% of its total duration. The full data-generation pipeline is illustrated in Figure 2.
We train SonicMaster by applying $19$ distinct degradations to the audio and pairing each with a matching editing instruction. The degradations span five classes: (i) EQ, (ii) Dynamics, (iii) Reverb, (iv) Amplitude, and (v) Stereo.
@@
Equalization (EQ): Spectral degradations cover $10$ effects: Brightness, Darkness, Airiness, Boominess, Muddiness, Warmth, Vocals, Clarity, Microphone, and X-band. Brightness, Darkness, Airiness, Boominess, and Warmth are emulated with low- or high-shelf EQ; Clarity with a Butterworth low-pass filter; Vocals and Muddiness with Chebyshev-II band-pass filters. Microphone applies one of $20$ Poliphone transfer functions (Salvi et al. 2025), while X-band uses an 881212-band, logarithmically spaced peaking EQ with $\textpm 6$ dB gain per band.
Dynamics: This category contains $2$ functions, Compression and Punch: modeled by feedforward compressor and transient shaper, respectively. Both these effects are applied in the time domain with a lossy, irreversible nature, as constructing an exact inverse of these is an ill-posed problem.
@@
Reverb: The Reverb category contains four functions. Three of them utilize the Pyroomacoustics library (Scheibler, Bezzam, and Dokmanić 2018), which simulates acoustic environments with the image source method. We simulate three types of rooms: Small, Big, and Mixed. For our fourth Reverb function, we utilize 12 selected room impulse responses from the openAIR library dataset (Howard and Angus n.d.), which give us audio with more real-life properties. The resulting impulse responses from all the functions are convoluted with the clean signals.
Amplitude: This category consists of two functions: Clipping and Volume. We clip the audio by raising the maximum amplitude to predefined levels, which causes distortion to arise. In Volume, we reduce the volume to a very low (barely audible) range, which decreases the signal-to-quantization noise ratio.
Stereo: This category contains one function to de-stereo the audio recording. First, the track is analyzed for suitability of this operation (whether it is ”stereo” enough) by taking the standard deviation of the difference betwen the left and the right channels and comparing to a set threshold of 0.08, then, if eligible, the left and right channels are combined to create a mono audio recording.
Each ground truth yields $7$ corrupted variants: $4$ with a single, $2$ with double, and $1$ with triple degradation. In multi-degradation, we sample at most one effect from each of the $5$ categories, so an EQ choice, for instance, blocks further EQ picks. To avoid duplicates in the single-degradation set, high-probability effects with narrow parameter ranges: Stereo, Clipping, and Punchare, are used only once.
Each degradation is linked to a one-sentence instruction randomly selected from 8–10 expert-written options; these sentences are concatenated into the full prompt, and we store two prompt variants per clip for robustness. We also record every applied effect and its parameters (gain, absorption), supporting tasks such as parameter prediction. For Compression and Reverb there is a 1515% chance of injecting “hidden clipping” with no corresponding instruction. When neither hidden clipping nor an Amplitude effect is present, the audio is peak-normalised to a random level between 0.81.00.8-1.0. Further details can be found in Appendix.
SonicMaster Architecture
SonicMaster combines Multimodal DiT (Esser et al. 2024) with subsequent DiT blocks (Peebles and Xie 2023). As outlined in Figure 3, the stereo waveform (44.1 kHz) ($x_{t}$) is first encoded by a VAE codec (Evans et al. 2024) into a compact spectro-temporal latent. Restoration therefore occurs entirely in this learned space, allowing large receptive fields without sample-level overhead. The MM-DiT block receives the degraded latent and the text embedding from a frozen FLAN-T5 encoder (Chung et al. 2024), and the conditioned representation is then passed to DiT layers that further process it to predict the velocity $v_{t}$ steering the latent toward its clean target $\hat{x}_{t}$. Prompts like ”reduce reverb” bias this prediction trajectory to suppress decay tails, while the downstream DiT layers also refine musical coherence. A pooled-audio branch, active in 2525% of training cases, concatenates a temporally averaged 5155–15 sec clean cue with the pooled prompt embedding and injects it at every MM-DiT/DiT layer, enabling seamless chaining of $30$ sec segments for long-form generation while degrading gracefully when no reference is supplied.
@@
We adopt the Stable Audio Open VAE (Evans et al. 2024) to encode–decode stereo signals sampled at 44.1 kHz, yielding a compact latent representation while retaining high-fidelity reconstruction. Text instructions are embedded with FLAN-T5 Large (Chung et al. 2024); the resulting tensor $c_{\text{text}}\in\mathbb{R}^{B\times S_{\text{text}}\times D_{\text{text}}}$ (with $D_{\text{text}}=1024$) is fed to the model as a conditioning signal.

### v2

@@ Abstract
Music recordings often suffer from audio quality issues such as excessive reverberation, distortion, clipping, tonal imbalances, and a narrowed stereo image, especially when created in non-professional settings without specialized equipment or expertise. These problems are typically corrected using separate specialized tools and manual adjustments. In this paper, we introduce SonicMaster, the first unified generative model for music restoration and mastering that addresses a broad spectrum of audio artifacts with text-based control. SonicMaster is conditioned on natural language instructions to apply targeted enhancements, or can operate in an automatic mode for general restoration. To train this model, we construct the SonicMaster dataset, a large dataset of paired degraded and high-quality tracks by simulating common degradation types with nineteen degradation functions belonging to five enhancements groups: equalization, dynamics, reverb, amplitude, and stereo. Our approach leverages a flow-matching generative training paradigm to learn an audio transformation that maps degraded inputs to their cleaned, mastered versions guided by text prompts. Objective audio quality metrics demonstrate that SonicMaster significantly improves sound quality across all artifact categories. Furthermore, subjective listening tests confirm that listeners prefer SonicMaster’s enhanced outputs over other baselines. The model and demo samples are available through https://github.com/AMAAI-Lab/SonicMaster.
1 Introduction
@@
Music recordings produced in amateur settings often suffer from a variety of quality issues that distinguish them from professionally mastered recordings (Wilson & Fazenda, 2016; Mourgela et al., 2024; Deruty & Tardieu, 2014). For instance, an enthusiast recording vocals in a garage may introduce excessive reverberation, making the voice sound distant and “echoey.” Similarly, using inexpensive microphones or misconfigured interfaces can lead to distortion and clipping when loud peaks exceed the recording range, resulting in harsh crackles or flattened dynamics (Zang et al., 2025). Tonal imbalances are also common: a home recording might sound overly “muddy” or “tinny” if certain frequency bands dominate or vanish due to poor room acoustics or improper microphone placement. Even the stereo image can be narrowed or skewed, reducing the sense of space in the mix. In practice, engineers address these problems with specialized tools: e.g., dereverberation plugins to remove room echo, declipping algorithms to reconstruct saturated peaks, equalizers to rebalance frequencies, and stereo enhancers to widen the image. Mastering a flawed track has become a labor-intensive process requiring expert skill and multiple stages of manual adjustment.
@@
The need for an automated all-in-one solution is evident. Creators with limited resources often lack the expertise to apply the right combination of restoration tools, and a piecemeal approach may fail to fully recover a track’s fidelity. This motivates SonicMaster, a unified approach to music restoration and mastering that can correct a broad spectrum of audio degradations within a single model. We introduce a single flow-based generative framework (Liu et al., 2022; Esser et al., 2024) that simultaneously performs dereverberation, equalization, declipping, dynamic-range expansion, and stereo enhancement. The backbone is trained on a curated corpus of polyphonic music rendered through a combinatorial grid of simulated degradations, enabling the network to learn the joint statistics and cross-couplings of common artifacts rather than treating them in isolation. This joint training eliminates the need for error-prone cascades of task-specific modules and reduces inference to a single forward pass.
Crucially, SonicMaster incorporates multimodal conditioning through natural language instructions that capture production objectives. A prompt such as reduce the hollow room sound attenuates late reflections without suppressing desirable early reverberation, whereas increase the brightness selectively enhances the treble frequencies while preserving spectral balance elsewhere.In the absence of a prompt, SonicMaster switches to an automatic mode that applies perceptually balanced mastering. Existing speech restoration models (e.g. VoiceFixer (Liu et al., 2021)) also address artifacts sequentially, ignoring their mutual influence. By unifying restoration and mastering tasks under a single, prompt-driven generative model, SonicMaster delivers professional-grade improvements while affording fine-grained creative control. Recent advances such as Mustango (Text-guided music generation) (Melechovsky et al., 2024), FlowSep (Yuan et al., 2025) (text-guided source separation), TangoFlux (Hung et al., 2024) (reward-optimized text-to-audio diffusion), and instruction-guided models like AUDIT (Wang et al., 2023) or AudioLDM/AudioLDM2 (Liu et al., 2023; 2024) illustrate powerful generative methods, but they target orthogonal tasks—generation, separation, or localized editing—rather than unified restoration. In contrast, SonicMaster uniquely addresses comprehensive multi-artifact music restoration and mastering through a single controllable rectified-flow architecture, bridging dereverberation, declipping, tonal rebalancing, dynamics, and stereo enhancement under prompt guidance.
In the absence of text-conditioned music-restoration data, we build a new large-scale corpus for controllable restoration. From $\approx 580$ $k$ Jamendo recordings, we retain $\approx 25$ $k$ high-quality 30-s segments, balanced across 10 genre groups by production quality score. Each clean clip is corrupted with one to three of 19 common effects drawn from five categories—EQ, dynamics, reverb, amplitude, and stereo—producing paired degraded versions. Every degraded sample is accompanied by a natural-language prompt describing the artifact or required fix, and all random effect parameters are stored as metadata. This genre-diverse collection of tens of thousands of prompt–audio pairs underpins SonicMaster training and offers a rigorous benchmark for controllable music-restoration research. Our main contributions are as follows:
  • We introduce SonicMaster, the first flow-matching model to simultaneously address $19$ common degradations, including reverb, EQ imbalance, clipping, dynamic range errors, and stereo artifacts in a single generative framework, eliminating sequential processing and cascading error.
  • Our SonicMaster enables precise user control through natural language conditioning, allowing targeted corrections (e.g., reduce hollow room sound for dereverberation) while maintaining autonomous operation when prompts are unavailable, bridging automated and user-directed restoration paradigms.
  • We construct and release [^1] https://github.com/AMAAI-Lab/SonicMasterthe first text-conditioned music-restoration corpus: 2525k high-fidelity Jamendo segments spanning 10 genres, each paired with $7$ degraded versions, detailed metadata, and a natural-language instruction describing the required fix, resulting in 175175k audio pairs.
2 Related Work
Restoring and mastering audio spans speech and music enhancement, audio inpainting, and source separation—areas that have mostly been handled separately (Záviška et al., 2020). Diffusion-based generative models and text-guided audio editing (Hou et al., 2025; Jiang et al., 2025; Zhang et al., 2024; Manor & Michaeli, 2024) now let us tackle these problems together. We review these advances, their uses, and the gaps that SonicMaster aims to fill. Early audio restoration efforts typically focused on single domains or isolated tasks, addressing issues like noise, clipping, or reverb in separation. Speech enhancement (Yousif & Mahmmod, 2025) and music enhancement evolved largely independently, and tasks such as audio inpainting or source separation were treated with specialized methods (Lemercier et al., 2025).
Audio Inpainting, Mixing and Declipping:
Early signal-model and interpolation methods could patch only very short gaps (<10<10 ms), leaving longer dropouts unresolved. Deep generative models now bridge that gap: diffusion-based systems convincingly regenerate missing music sections and clipped peaks (Moliner & Välimäki, 2023). The authors in (Wang et al., 2023; Liu et al., 2023) extend this with instruction-guided diffusion for audio inpainting, while VoiceFixer (Liu et al., 2021) jointly denoises, dereverbs, and declipse speech, though it is restricted to voice and does not offer user control. In music, (Imort et al., 2022) removed heavy guitar distortion (including clipping) with neural networks, surpassing sparse-optimization baselines in quality and speed. (Lee et al., 2024) introduce a pruning approach to recover sparse audio effect chains from mixed recordings, essentially reverse-engineering mixing graphs from input/output pairs. Alongside works like (Bhandari et al., 2025) on iterative corruption refinement, (Steinmetz et al., 2021) on differentiable mixing consoles and (Martínez-Ramírez et al., 2022) on out-of-domain mixing generalization, and diffusion restorers such as MaskSR (Li et al., 2024), these highlight emerging methods that bridge audio restoration with controllable, interpretable effect modeling.
Equalization and Tonal Restoration:
Research on learning-based equalization is still emerging. In recently proposed work (Mockenhaupt et al., 2024), the authors introduce CNN-based approach to automatically equalize instrument stems by predicting parametric EQ settings, showing improvements over earlier heuristic methods. Notably, the VoiceFixer (Liu et al., 2021) addressed bandwidth extension, essentially restoring high-frequency content as part of its speech restoration, which can be seen as a form of equalization correction. Similarly, diffusion-based restorers like MaskSR (Li et al., 2024) treat low-frequency muffling as a distortion to fix, using discrete token prediction to restore a balanced spectrum. In Text2FX (Chu et al., 2025), CLAP (Elizalde et al., 2023) is used in inference mode to steer the parameters of EQ and reverb audio effects.
3 Method
3.1 Dataset
In the absence of text-conditioned music restoration datasets, we generate SonicMaster dataset by pairing high-quality audio with systematically applied degradations and corresponding natural language instructions. Our source comprises songs from (Roy et al., 2025) and additional content from Jamendo [^2] https://www.jamendo.com/ under Creative Commons licence using the official Jamendo API. In total, we have sourced 580580k recordings. We ensure balanced genre representation by defining $10$ groups, where each group consists of multiple semantically related genre tags, e.g., Hip-Hop genre group containing the following tags: ”rap”, ”hiphop”, ”trap”, ”alternativehiphop”, ”gangstarap”. Complete taxonomies are provided in the Appendix. Track selection employs Audiobox Aesthetics toolbox (Tjandra et al., 2025) for automated production quality assessment. We select $2,500$ songs per genre groups using adaptive production quality thresholds ranging from $6.5$ to $8$ to balance comprehensive sub-genre representation with sufficient production quality. We extract random 30s excerpts from each track, positioned between 1515% and 8585% of its total duration. The complete pipeline is illustrated in Figure 1. We train SonicMaster by applying $19$ distinct degradations to the audio and pairing each with a matching natural language editing instruction. The degradations span five classes: (i) EQ, (ii) Dynamics, (iii) Reverb, (iv) Amplitude, and (v) Stereo.
@@ 3.1 Dataset
Equalization (EQ): Spectral degradations cover $10$ effects targeting perceptual audio characteristics: Brightness, Darkness, Airiness, Boominess, Muddiness, Warmth, Vocals, Clarity, Microphone, and X-band. Brightness, Darkness, Airiness, Boominess, and Warmth are emulated with low- or high-shelf EQ; Clarity with a Butterworth low-pass filter; Vocals and Muddiness with Chebyshev-II band-pass filters. Microphone applies one of $20$ Poliphone transfer functions (Salvi et al., 2025), while X-band uses an 881212-band, logarithmically spaced peaking EQ with $\textpm 6$ dB gain per band.
Dynamics: Temporal envelope modification via two functions: Compression (feedforward dynamic range compression) and Punch (transient shaping). Both exhibit lossy, non-invertible characteristics, rendering exact restoration mathematically ill-posed and requiring learned approximations.
Reverb: The Reverb category contains four distinct approaches: three of them utilise the Pyroomacoustics library (Scheibler et al., 2018), which simulates acoustic environments with the image source method. We simulate three types of rooms: Small, Big, and Mixed. For our fourth Reverb function, we utilize 12 selected room impulse responses from the openAIR library dataset (Howard & Angus, n.d.), which give us audio with more real-life properties. The resulting impulse responses from all the functions are convoluted with the clean signals.
Amplitude: Two complementary degradations target signal amplitude: Clipping/Volume. Clipping introduces hard nonlinear distortion by constraining peak amplitudes to predefined thresholds; Volume reduction attenuates signals to near-inaudible levels, degrading the signal-to-quantization-noise ratio and simulating poor recording practices.
Stereo: A function to de-stereo the audio recording – tracks undergo stereo content analysis via left-right channel difference standard deviation (threshold: 0.08); qualifying recordings are converted to monophonic by channel summation, simulating poor mixing or playback equipment limitations.
Each ground truth yields $7$ corrupted variants: $4$ with a single, $2$ with double, and $1$ with triple degradation. In multi-degradation, we sample at most one effect from each of the $5$ categories, so an EQ choice, for instance, blocks further EQ picks. To avoid duplicates in the single-degradation set, high-probability effects with narrow parameter ranges: Stereo, Clipping, and Punchare, are used only once. Each degradation is linked to a one-sentence instruction randomly selected from 8–10 expert-written options; these sentences are concatenated into the full prompt, and we store two prompt variants per clip for robustness. We also record every applied effect and its parameters (gain, absorption), supporting tasks such as parameter prediction. For Compression and Reverb, there is a 1515% chance of injecting “hidden clipping” with no corresponding instruction. When neither hidden clipping nor an Amplitude effect is present, the audio is peak-normalised to a random level between 0.81.00.8-1.0. Further details can be found in Appendix.
3.2 SonicMaster Architecture
SonicMaster employs a hybrid architecture combining Multimodal Diffusion Transformer (MM-DiT)(Esser et al., 2024) blocks with subsequent Diffusion Transformers (DiT) layers (Peebles & Xie, 2023). As outlined in Figure 2, stereo waveforms (44.1 kHz) ($x_{t}$) undergo VAE ecoding (Evans et al., 2024) into compact spectro-temporal latent representations. Restoration, therefore, occurs entirely in this learned space, allowing large receptive fields without sample-level overhead. The MM-DiT processes degraded latent representations alongside the text embeddings from a frozen FLAN-T5 encoder (Chung et al., 2024). The resulting conditioned representations pass through subsequent DiT layers to predict flow velocity $v_{t}$, steering the latent toward its clean target $\hat{x}_{t}$. Prompts like ”reduce reverb” biases this prediction trajectory to suppress decay tails, while the downstream DiT layersrefine musical coherence. A pooled-audio branch, active in 2525% of training cases, concatenates a temporally averaged 5155–15s clean cue with the pooled prompt embedding and injects it at every MM-DiT/DiT layer, enabling seamless chaining of $30$s segments for long-form generation while degrading gracefully when no reference is supplied.
@@ 3.2 SonicMaster Architecture
We adopt the Stable Audio Open VAE (Evans et al., 2024) to encode–decode stereo signals sampled at 44.1 kHz, yielding a compact latent representation while retaining high-fidelity reconstruction. Text instructions are embedded with FLAN-T5 Large (Chung et al., 2024); the resulting tensor $c_{\text{text}}\in\mathbb{R}^{B\times S_{\text{text}}\times D_{\text{text}}}$ (with $D_{\text{text}}=1024$) is used as a conditioning signal.
Rectified Flow Training:
SonicMaster employs rectified flow (Liu et al., 2022; Esser et al., 2024), to predict flow velocity from degraded to clean audio in latent space, unlike other models that map noise to output distributions (Fei et al., 2024; Hung et al., 2024).
We assign timestep $t=1$ to the latent representation of the degraded audio $x_{1}$, and $t=0$ to the latent representation of the clean audio target $x_{0}$. During training, we feed the model with samples $x_{t}$, which are linear interpolations between degraded input $x_{1}$ and clean target $x_{0}$:
$x_{t}=tx_{1}+(1-t)x_{0}$ (1)
where timestep $t$ is drawn from a skewed distribution $p(t)=0.5U(t)+t,\quad t\in[0,1]$ with increasing probability for higher $t$, where $U$ represents a uniform distribution. This skewed distribution gives emphasis to more degraded inputs given the interpolation of training data in Eq. 1. The model is trained to predict the flow velocity $v_{t}$ from the current $x_{t}$ to the target clean audio $x_{0}$: $v_{t}=-\frac{dx_{t}}{dt}=x_{0}-x_{1}$. The model $f_{\theta}$ with parameters $\theta$ estimates the velocity $\hat{v_{t}}$, $f_{\theta}(x_{t},t,c_{text})=\hat{v_{t}}$, where $c_{text}$ is the text condition from the FLAN-T5 model, which is passed to the dual-stream MM-DiT blocks as one of the streams. The $c_{text}$ condition is also passed through a pooled projection and used to control the scale and shift factors of the adaptive layer-norm layers in both MM-DiT and DiT blocks. The training loss is then given as:
[…]
Inference transforms degraded audio input $x_{1}$ to clean audio output $x_{0}$ by integrating the predicted velocity $\hat{v_{t}}$ using the forward Euler method: $x_{t-h}=x_{t}+h\hat{v_{t}}$, where $h\in[0,1]$ is the step computed as the inverse of the total timesteps dedicated for integration.
Inference: During inference, SonicMaster takes in an audio input and a text instruction given by the user to perform the desired restoration/mastering operation. Inference is possible without text input in the so-called auto-correction mode. To process full-length songs, SonicMaster operates on chunks of 30s and then connects the segments together. After the first segment is inferred, the last 10s of this output are used to condition the next segment inference through the audio pooling branch. The overlapping regions of the resulting segments are then linearly interpolated over the overlapping 10s to connect the segments together.
4 Experimental Setup and Baselines
4.1 Baselines and training setup
We train SonicMaster using 5 NVIDIA L40S GPUs for 40 epochs with a total batch size of 80. We adopt classifier-free guidance (Ho & Salimans, 2022) by (i) dropping the text prompt in 1010% of samples and (ii) replacing it in another 1010% with one of four generic phrases (“Make it sound better!”, “Master this track for me, please!”, “Improve this!”, “Can you improve the sound of this song?”). In 2525% of cases, the model is additionally conditioned—via the pooling branch—on the first 10 s of clean audio. Unless stated otherwise, all experiments follow these conditioning settings while comparing multiple SonicMaster variants and baselines.
[…]